HETAV PATEL

Grade 11,

find the pressure of 16g of O2 gas in the atm present in a flask of 1.12 lit at 273K

find the pressure of 16g of O2 gas in the atm present in a flask of 1.12 lit at 273K

Grade:11

1 Answers

Vikas TU
7 years ago
From the ideal gas eqn.,
PV =nRT
P*1.12 = (16/32)*0.0821*273
P = > 11.2066/1.12 = > 10.005 atm.
This would b the final pressure  in the flask.